OpenStudy (anonymous):

Solve for x: 2e^6x = 18

OpenStudy (campbell_st):

divide both sides by 2 \[e^{6x} = 9\] you will use log laws now in particular the log of the exponential \[\log _{a}a^x = x\] take the base e log of both sides \[\ln(e^{6x}) = \ln(9)\] this should be enough to help you with a solution